Biologists want to set up a station to test alligators in the lake for West Nile Virus. Suppose that the cost for such a station are $2,500 for setup costs and $3.00 to administer each test.
a. Write an expression that gives the total cost to test x animals.
b. You can find the average cost per animal by dividing total costs by number of animals. Write the expression that gives the average cost per animal.
c.Find the average cost per animal for 10 animals, 100 animals, and 1,000 animals.
d.As the number of animals increases, what happens to the average cost to test the animals? Would the average cost ever fall below $3.00? If so, identify a value that supports your answer.If not explain how you know.
e. How many animals should be testes for the average cost to be $5.00 per animal?

Answer by stanbon(75887)

Biologists want to set up a station to test alligators in the lake for West Nile Virus. Suppose that the cost for such a station are $2,500 for setup costs and $3.00 to administer each test.
a. Write an expression that gives the total cost to test x animals.
C(x) = 2500 + 3x
------------------------------
b. You can find the average cost per animal by dividing total costs by number of animals. Write the expression that gives the average cost per animal.
Ave = (2500+3x)/x
------------------------------
c.Find the average cost per animal for 10 animals, 100 animals, and 1,000 animals.
Ave(10) = (2530)/10 = $253
Ave(100) = (2800)/100 = $28
Ave(1000) = (5500)/1000 = $5.50
------------------------------
d.As the number of animals increases, what happens to the average cost to test the animals? Would the average cost ever fall below $3.00? If so, identify a value that supports your answer.If not explain how you know.
---
Horizontal asymptotte at Ave(x) = 3
===========================================
e. How many animals should be tested for the average cost to be $5.00 per animal?
Solve: 5 = (2500+3x)/x
5x = 2500+3x
2x = 2500
x = 1250
===============
Cheers,
Stan H.
================